An assembly line has three operations in sequence.  The work-centers take 5, 3, and 2 minutes per job. (So job 1 enters the first work-center, where it takes 5 minutes; then it is moved to the second work center where it takes 3 minutes while job 2 enters the first work-center, etc.)

In the above line, if the second work-center operation time is reduced by ten percent, what is the throughput of this line (as number of jobs completed per hour in the long run)?

(Provide answer to two decimal places)
Answer #1

In the initial case, we can see that work center take 5,3 and 2 minutes respectively. Hence first work center's hourly capacity = 60/5 = 12. Similarly second and third work center's capacity is 60/3 = 20 and 60/2 = 30 per hour.

As we can see the bottleneck for the line is work center 1 as it has the lowest capacity. Hence the current throughput in terms of number of jobs completed per hour in the long run is 12 per hour. Now if we reduce the operation time of second work center by 10 percent, the bottleneck station is still work center 1. Hence the throughput of the line remains the same as before which is 12 jobs per hour.